# Explaining Sources of Uncertainty in Automated Fact-Checking

### v1

@@ H.1 Participants and Materials
We recruited N=12 participants from Prolific, screened to be native English speakers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the United States. The study was approved by our institution’s Research Ethics Committee (reference number 504-0516/24-5000).
@@
Forty instances (20 from DRUID, 20 from HealthVer) were selected at random for evaluation. For each instance, participants were provided with a claim, two evidence documents, model verdict, model numerical certainty, and three alternative explanations (see Figure 6 in H.6). The explanations presented to participants were those generated using Qwen2.5-14b-instruct Qwen Team (2024) based on its automatic evaluation performance. Each participant evaluated explanations for 10 instances (5 labelled ‘True’, 5 labelled ‘False’), in addition to two attention check instances which were used to screen responses for quality.
@@
Participants read information about the study (see H.3) and provided informed consent (see H.4) before reading detailed task instructions and completing a practice example of the task (see H.5). Participants then progressed through the study at their own pace. The task took approximately 20 minutues, and participants were paid £3 for their work.
H.2 Human Evaluation Results
H.2.1 Interrater agreement
@@
H.3 Human Evaluation Information Screen
@@ 1. What is the project about?
Our goal is to make sure that AI fact-checking systems can explain the decisions they produce in ways that are understandable and useful to people. This survey is part of a project to help us understand what kinds of explanations are helpful and why.
@@ 3. Source of funding
This project has received funding from the ERC (European Research Council) Starting Grant on Explainable and Robust Fact Checking under grant agreement ID no. 101077481.
@@ 7. Participants’ rights under the General Data Protection Regulation (GDPR)
As a participant in a research project, you have a number of rights under the GDPR. Your rights are specified in the University of Copenhagen’s privacy policy. https://informationssikkerhed.ku.dk/english/protection-of-information-privacy/privacy-policy/
@@ 8. Person responsible for storing and processing of data
University of Copenhagen, CVR no. 29979812, is the data controller responsible for processing data in the research project.
The research project is headed by Prof. Isabelle Augenstein who can be contacted via email: augenstein@di.ku.dk, phone: <>, address: Øster Voldgade 3 1350 Copenhagen, Denmark.
Greta Warren is the contact point for this project and can be contacted via email: grwa@di.ku.dk, phone: <>, address: Øster Voldgade 3, 1350 Copenhagen, Denmark.
@@
H.4 Human Evaluation Consent Form
@@
  • I hereby give my consent that the University of Copenhagen may register and process my data as part of the Human-Centred Explainable Fact Checking project.
@@
H.5 Evaluation Task Instructions
@@
In this study you will see claims, an AI system’s prediction about whether this claim is true or false, how certain the system is about its label, and the corresponding evidence used to make the prediction. You will also see three different explanations for why the AI system is certain or uncertain about its prediction. These explanations are intended help you decide how to interpret the true/false prediction.
@@
Important: Please only consider the provided information (claim, evidence documents, and explanations) when evaluating explanations. Sometimes you will be familiar with the claim, but we ask you to approach each claim as new, whether or not you have seen it before. It doesn’t matter whether you personally agree or disagree with the claim or evidence – we are asking you to evaluate what the AI produces: if you were to see this claim for the first time, would you find the explanation provided by the AI useful? On the next page, you will see an example of the task.
H.6 Example of human evaluation set-up

### v2

@@ I.1 Participants and Materials
We recruited N=12 participants from Prolific, screened to be native English speakers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the United States. The study was approved by our institution’s Research Ethics Committee (reference number anonymised).
@@
Forty claims (20 from DRUID, 20 from HealthVer) were selected at random for evaluation, while 120 explanations were evaluated in total. For each instance, participants were provided with a claim, two evidence documents, model verdict, model numerical certainty, and three alternative explanations (see Figure 6 in I.6). The explanations presented to participants were those generated using Qwen2.5-14b-instruct Qwen Team (2024) based on its automatic evaluation performance.
Each participant evaluated explanations for 10 instances (5 labelled ‘True’, 5 labelled ‘False’), in addition to two attention check instances which were used to screen responses for quality. To mitigate the difficulty and cognitive demands of the task, we kept the number of instances per participant low, and removed participants who failed attention checks from the data. In order to minimise the risk of bias, each explanation presented to participants was assigned a neutral label: ‘Explanation A’, ‘Explanation B’, ‘Explanation C’, so that participants could not infer anything about how the explanations were generated or indicate that one should be favoured over another. Each explanation was also presented in a different, random order for each instance (e.g., ABC for Claim 1, CAB for Claim 2, etc.) to mitigate order effects such as primacy bias or recency bias. The order in which the response options ‘Explanation A’, ‘Explanation B’, ‘Explanation C’ also appeared in the same random order.
@@
Participants read information about the study (see I.3) and provided informed consent (see I.4) before reading detailed task instructions and completing a practice example of the task (see I.5). Participants then progressed through the study at their own pace. The task took approximately 20 minutes, and participants were paid £3 for their work.
I.2 Human Evaluation Results
I.2.1 Interrater agreement
@@
I.2.2 CLUE Variant preferences
PromptBase CLUE-S CLUE-Steer
Helpfulness
Overall 2.025 1.892 1.867
DRUID 1.9 1.917 1.767
HealthVer 2.15 1.867 1.967
Consistency
Overall 1.875 1.783 1.817
DRUID 1.717 1.75 1.617
HealthVer 2.033 1.817 2.017
Non-redundancy
Overall 2.05 1.908 1.833
DRUID 1.983 1.983 1.683
HealthVer 2.117 1.833 1.983
Coverage
Overall 1.967 1.775 1.758
DRUID 1.767 1.75 1.617
HealthVer 2.167 1.8 1.9
Overall Quality
Overall 1.967 1.908 1.925
DRUID 1.9 1.9 1.817
HealthVer 2.033 1.917 2.033
Table 11: Mean Average Rank (MAR) for the five human-evaluation criteria applied to explanations from Qwen2.5-14B-Instruct on the HealthVer and DRUID datasets (chosen for its high faithfulness; see §4.5). PromptBaseline, CLUE-Span (CLUE-S), and CLUE-Span+Steering (CLUE-Steer) are compared. Lower MAR means a better (higher) average rank; the best score in each row is boldfaced.
Table 11 shows the mean rank assigned by participants to each explanation variant. The explanations generated by CLUE were preferred by our participants to those generated using PromptBaseline: the explanations generated by CLUE-Span+Steering were rated as most helpful, highest coverage, and containing the least amount of redundant information, while those from CLUE-Span were judged to have the highest consistency and overall quality.
I.3 Human Evaluation Information Screen
@@ 1. What is the project about?
Our goal is to make sure that AI fact-checking systems can explain the decisions they produce in ways that are understandable and useful to people. This survey is part of a project to help us understand [anonymised]
@@ 3. Source of funding
This project has received funding from [anonymised]
@@ 7. Participants’ rights under the General Data Protection Regulation (GDPR)
As a participant in a research project, you have a number of rights under the GDPR. Your rights are specified in [anonymised]
@@ 8. Person responsible for storing and processing of data
[anonymised]
@@
I.4 Human Evaluation Consent Form
@@
  • I hereby give my consent that [anonymised] may register and process my data
@@
I.5 Evaluation Task Instructions
@@
In this study you will see claims, an AI system’s prediction about whether this claim is true or false, how certain the system is about its label, and the corresponding evidence used to make the prediction. You will also see three different explanations for why the AI system is certain or uncertain about its prediction. These explanations are intended to help you decide how to interpret the true/false prediction.
@@
Important: Please only consider the provided information (claim, evidence documents, and explanations) when evaluating explanations. Sometimes you will be familiar with the claim, but we ask you to approach each claim as new, whether or not you have seen it before. It doesn’t matter whether you personally agree or disagree with the claim or evidence - we are asking you to evaluate what the AI produces: if you were to see this claim for the first time, would you find the explanation provided by the AI useful? On the next page, you will see an example of the task.
I.6 Example of human evaluation set-up